Subject: Thin films mornng report 4/30/04
From: james haydon <jhaydon@snf.stanford.edu>
Date: Fri, 30 Apr 2004 09:37:00 -0700
Fri, 30 Apr 2004 09:37:00 -0700
           

Thinfilms Morning Report 4/30/2004

 

 

Gryphon:            ok

Innotec;            ok

Balzers :            ok

STS Dep:            ok

Implanter:            phosphine bottle change

Metalica:            cryo pump replacement 5/3/04

SCT:                completion of installation near end of May

CPD:                ok                   

WbGen-CTB:            control temp bath not working should be back up 
late next week

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            right hot pot continues to drain after drain has 
been turned off

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 



           

Thinfilms Morning Report 4/30/2004

 

 

Gryphon:            ok

Innotec;            ok

Balzers :            ok

STS Dep:            ok

Implanter:            phosphine bottle change

Metalica:            cryo pump replacement 5/3/04

SCT:                completion of installation near end of May

CPD:                ok                   

WbGen-CTB:            control temp bath not working should be back up late next week

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            right hot pot continues to drain after drain has been turned off

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok